OpenStudy (anonymous):

15. A recursive function is shown below: f(1) = 2 and f(n) = f(n – 1) – 5; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? (4 points) A. 2, 3, 8, 13, 18, ... B. 2, 7, 12, 17, 22, ... C. 2, –3, –8, –13, –18, ... D. 2, –7, –12, –17, –22, ...

OpenStudy (amoodarya):

put n=2,3,4,... to find that like this f(2)=f(1)-5=2-5=-3 f(3)=f(2)-5 = -3-5=-8

OpenStudy (amoodarya):

it is arithmetic progression

OpenStudy (anonymous):

Ok thanks